OpenStudy (anonymous):

parametric equations...please help me with last 3...i am totally confused...explain them

OpenStudy (amistre64):

arclength is just\[\int ds=\int_{a}^{b}\sqrt{(x')^2+(y')^2}~dt\]

OpenStudy (anonymous):

i know tht but how to place the functions

OpenStudy (amistre64):

v in terms of s? i assume that they are refering to s(t) as a position equation maybe? then ds = v(t)

OpenStudy (amistre64):

replace v by ds/dt \[\frac12m(\frac{ds}{dt})^2+mgy=k\]

OpenStudy (amistre64):

solve for dt

OpenStudy (anonymous):

\[dt = \frac{ ds }{ \sqrt{1+2g(1-y)} }\]

OpenStudy (amistre64):

k = mg \[\frac12m(\frac{ds}{dt})^2+mgy=mg\] \[\frac12(\frac{ds}{dt})^2+gy=g\] \[(\frac{ds}{dt})^2=2(g-gy)\] \[\frac{ds}{\sqrt{2g(1-y)}}=dt\]

OpenStudy (anonymous):

now how to express the arc length

OpenStudy (amistre64):

it says to express it in terms of f' and g' wrt u \[\int_{}^{}\sqrt{(\frac{df}{du}\frac{du}{dt})^2+(\frac{dg}{du}\frac{du}{dt})^2}~dt\] \[\int_{}^{}\sqrt{(\frac{du}{dt})^2\left((\frac{df}{du})^2+(\frac{dg}{du})^2\right)}~dt\] \[\int_{}^{}\sqrt{(\frac{df}{du})^2+(\frac{dg}{du})^2}~\frac{du}{dt}dt\] \[\int_{0}^{\pi}\sqrt{(f')^2+(g')^2}~du\]

OpenStudy (amistre64):

pfft, f and h, but samesame

OpenStudy (amistre64):

the nest thing is a straight line equation between 2 points .....

OpenStudy (anonymous):

how did u do that lol

OpenStudy (amistre64):

*next

OpenStudy (anonymous):

so that's the answer for f)

OpenStudy (amistre64):

thats what i get for "f" yes

OpenStudy (amistre64):

but i used a g instead of an h :/

OpenStudy (anonymous):

g stands for gravity so should i just replace g by h?

OpenStudy (amistre64):

since y = h(u), yes .... i simply misread, mistyped, it is all

OpenStudy (anonymous):

haha thanks...how to do g?

OpenStudy (amistre64):

if i asked you to find the equation of a line between say: (0,1) and (5,0) what would you do?

OpenStudy (anonymous):

y-y0 = m(x-x0)

OpenStudy (amistre64):

yep, thats all g is asking for ... and also to put it into parametric form

OpenStudy (anonymous):

so y=-0.2x+1 ????

OpenStudy (amistre64):

first determine the cartesian, xy plane, y=mx+b format

OpenStudy (amistre64):

good, and with that we can see that: x=x y=mx+b

OpenStudy (amistre64):

just use the points they gave you instead

OpenStudy (anonymous):

\[y = -\frac{ 2 }{ \pi }x+1\]

OpenStudy (anonymous):

what next ?

OpenStudy (amistre64):

thats the cartesian form the parametric form is: x = x y = -2/pi x + 1 but replace x on the right by t

OpenStudy (anonymous):

ok lemme try that

OpenStudy (anonymous):

oh just replace x by t???

OpenStudy (amistre64):

yeah

OpenStudy (anonymous):

y = -2/pi t + 1???

OpenStudy (amistre64):

x = t y = -2t/pi + 1 this gives us parametric equations in terms of "t", but i have to be leery as to if this is "correct" to do or not

OpenStudy (anonymous):

i thought parametric equations is something like x^2 + y^2 = r^2 where r is radius...and this is cartesian equation

OpenStudy (amistre64):

nah, thats polar

OpenStudy (amistre64):

parametric simply define x and y seperately

OpenStudy (anonymous):

oh then what is cartesian

OpenStudy (amistre64):

y = f(x) is cartesian ... parametric define x and y seperately by parameters x = f(t) y = g(t)

OpenStudy (anonymous):

ohk...so cartesian would be f(t) = -2t/pi + 1

OpenStudy (amistre64):

we can define a vector setup in terms of:\[\vec r(t)=(x(t),y(t))\]

OpenStudy (amistre64):

no, cartesian is the usual xy plane .. y = f(x) = mx+b type stuff, defining y in terms of x

OpenStudy (anonymous):

f(x) = -(2/pi)x + 1 ??

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

and for time just replace y by zero and solve for t...right?

OpenStudy (amistre64):

at t=0, x=0 at t=0, y=1 at t=?, x=pi/2 at t=?, y=0

OpenStudy (amistre64):

im just not sure if its a good or proper idea to equate x=t .... the speed determines the time it takes to travel the slope, not the shape of the slope

OpenStudy (anonymous):

x=t so at x = pi/2, t=pi/2 at y=0, t= pi/2 again

OpenStudy (amistre64):

it says at the start that x and y are functions of u so lets use u instead of t

OpenStudy (anonymous):

oh man where does this u come from

OpenStudy (amistre64):

3rd sentence

OpenStudy (amistre64):

when u=0, x=0 when u=pi, x=pi/2 let x = u/2

OpenStudy (amistre64):

\[x = x\\ y = -\frac{2}{\pi}x+1 \] \[x = f(u) = \frac u2\\ y = h(u)=-\frac{2}{\pi}\frac u2+1=-\frac u{\pi}+1 \]

OpenStudy (amistre64):

\[f' = \frac12\\h' = -\frac{1}{\pi}\] \[\int_{u=0}^{u=pi}~\sqrt{(f')^2+(h')^2}du\] \[\int_{u=0}^{u=pi}~\sqrt{\frac14+\frac{1}{\pi^2}}~du\]

OpenStudy (anonymous):

1.86

OpenStudy (amistre64):

\[\int_{u=0}^{u=pi}~\frac{\sqrt{\pi^2+4}}{2\pi}~du\]

OpenStudy (amistre64):

1.86 yes, but not sure how that would relate to time

OpenStudy (anonymous):

18.38

OpenStudy (anonymous):

wait y do we got two different equations

OpenStudy (amistre64):

we dont ... one is a simplification is all

OpenStudy (anonymous):

i got different answer for second one

OpenStudy (amistre64):

sqrt(pi^2+4)/2

OpenStudy (anonymous):

5.85

OpenStudy (anonymous):

wait without integral it is 1.86

OpenStudy (anonymous):

ya i integrate it...sry abt that

OpenStudy (anonymous):

lemme check our answers with @electrokid

OpenStudy (amistre64):

how does this relate u to t tho?

OpenStudy (amistre64):

im thinking that line "e" is important

OpenStudy (anonymous):

@electrokid already helped me solve that...we went upto there then i got confused and couldnt solve anything...thx for ur help...i just need him to verify answers

OpenStudy (amistre64):

k